how many ounces of 20% hydrochloric acid solution and 50% hydrochloric acid solution must be mixed together to get 12 ounces of 30% hydrochloric acid solution
Answer by Boreal(15235)

x=ounces of 20%
12-x=ounces of 50%
0.20x + 0.50(12-x) =0.3acid*12ounces
0.20x + 6-0.50x =3.6 acid ounces
-0.30x=-2.4; divide both sides by -0.30
x=8 ounces ;; 20% ;; 8*0.20=1.6
12-x=4 ounces ;;50% ;; 4*0.50=2
Sum is 3.6 acid-ounces